use the method of substitution to solve the system of leanear equations
2x-y=8
7x-3y=28
what is the solution?

Answer by TimothyLamb(4379)

2x - y = 8
7x - 3y = 28
---
2x - y = 8
2x - 8 = y
---
7x - 3y = 28
7x - 3(2x - 8) = 28
7x - 6x + 24 = 28
---
x = 4
y = 0
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php